OpenStudy (anonymous):

After three days of hiking, a group of hikers are 30 km north from their starting position. On the first day, they hiked 20 km east; on the second day, they hiked 30 km 53degrees North of West. Calculate their displacement on the third day. Show solution.

OpenStudy (binary3i):

 0-1=30 1-2=20 2-3=30 smaller angle between 1-2 and 2-3 is 53 just find the distance 0-3 using pythagoras